OpenStudy (kingj):

If g(x) = 3x + 5, what is g^–1(x)?

OpenStudy (anonymous):

To find the inverse, you need to "undo" the actions of the original function. eg, g(4) = 3*4 + 5 = 17. How do you get from 17 back to 4? first, you would want to subtract 5, leaving 12. Then you can divide by 3. Thus, \[g^{-1}(x) = 1/3(x-5)\]

OpenStudy (precal):

g(x)=3x+5 rewrite as y=3x+5 step 1: switch x and y x=3y+5 Step 2: solve for y x=3y+5 x-5=3y (x-5)/3=y (1/3)x-(5/3)=y^-1

OpenStudy (precal):

Note that both answers given are the same, just different variation